# Enhancing Object Detection Accuracy in Underwater Sonar Images through Deep Learning-based Denoising

### v1

@@ Abstract
Sonar image object detection is crucial for underwater robotics and other applications. However, various types of noise in sonar images can affect the accuracy of object detection. Denoising, as a critical preprocessing step, aims to remove noise while retaining useful information to improve detection accuracy. Although deep learning-based denoising algorithms perform well on optical images, their application to underwater sonar images remains underexplored.
This paper systematically evaluates the effectiveness of several deep learning-based denoising algorithms, originally designed for optical images, in the context of underwater sonar image object detection. We apply nine trained denoising models to images from five open-source sonar datasets, each processing different types of noise. We then test the denoised images using four object detection algorithms. The results show that different denoising models have varying effects on detection performance. By combining the strengths of multiple denoising models, the detection results can be optimized, thus more effectively suppressing noise.
Additionally, we adopt a multi-frame denoising technique, using different outputs generated by multiple denoising models as multiple frames of the same scene for further processing to enhance detection accuracy. This method, originally designed for optical images, leverages complementary noise-reduction effects. Experimental results show that denoised sonar images improve the performance of object detection algorithms compared to the original sonar images.
Index Terms:
Image Denoising, Underwater Sonar Images, Object Detection, Multi-frame DenoisingI Introduction
Underwater sonar imaging plays an indispensable role in marine exploration and various ocean industries, providing valuable insights into underwater environments. Unlike optical imaging, where light propagation is restricted, sonar systems utilize sound waves that travel farther, allowing them to cover larger underwater areas. This makes sonar images an ideal choice for applications such as seabed mapping, underwater object detection, and navigation. However, despite the advantages of sonar imaging, its image quality is often severely compromised by noise, which negatively impacts the accuracy of downstream tasks, such as object detection.
In sonar images, noise can originate from various factors, including environmental interference, sensor imperfections, and the inherent characteristics of sound wave propagation in water. Common types of sonar image noise include Gaussian noise, speckle noise, and Poisson noise. Gaussian noise typically arises from random fluctuations in sensor readings or environmental changes. Speckle noise, caused by sound wave scattering, manifests as granular interference, which can obscure object boundaries. Poisson noise, a type of noise induced by the Poisson distribution in statistics, primarily stems from the natural variability of the signal, particularly when the signal generation process is random. These noise types severely degrade the signal-to-noise ratio (SNR) of sonar images, resulting in poor image visibility and consequently affecting the performance of object detection tasks.
A standard approach for improving the quality of sonar images and enhancing object detection accuracy is through denoising. Denoising algorithms aim to reduce or eliminate noise while preserving critical features in the image. Traditional denoising methods, such as spatial filtering and thresholding, often perform poorly in underwater sonar image processing due to the complexity of the noise. These methods may blur or distort important details, such as object contours and features, which are vital for accurate detection. As a result, an increasing number of studies have turned to more advanced techniques, particularly deep learning-based methods, which can adapt to the unique characteristics of underwater sonar images.
Deep learning-based denoising technologies, especially self-supervised learning models, have demonstrated significant potential in handling image noise. These algorithms do not rely on large amounts of labeled data but instead learn to denoise by utilizing the inherent structure of the data itself. These models are particularly well-suited for scenarios where clear ground-truth images are unavailable, such as in underwater sonar imaging. In addition to directly denoising the original images, another promising approach is to treat multiframe denoised images as different exposures of the same scene and apply multi-frame image denoising techniques to further reduce noise. This technique integrates information from multiframe denoised images, thereby reducing noise while preserving more details and structural information. Multi-frame image denoising has been widely studied and applied in the field of optical images, demonstrating significant advantages in improving image quality. Although this method has been relatively less explored in underwater sonar images, its potential in underwater environments is equally noteworthy, especially when dealing with sonar images contaminated by noise.
This paper proposes an integrated approach to systematically evaluate the impact of denoising and multi-frame image denoising strategies on underwater sonar image object detection performance. We not only focus on how denoising techniques can improve image quality but also explore the application of multi-frame image denoising and assess its effectiveness in enhancing object detection accuracy and robustness.
We conducted experiments on several underwater datasets, including FDD[39], UATD[38], SCTD[40], MarineDebris[41], and 2020Contest[42]. These datasets feature diverse underwater scenes and various target types, enabling an effective evaluation of the performance of denoising, object detection, and image fusion methods . We combined several advanced deep learning techniques, employing multiple denoising algorithms and object detection models to assess their performance on underwater sonar images. Specifically, we used three self-supervised denoising algorithms Neighbor2Neighbor[2], Blind2Unblind[5], and DSPNet[6], aimed at removing noise from sonar images and restoring critical image information. For object detection, we evaluated four different models: SSD-MobileNetV2[27][47], Faster R-CNN[31], SSD300[27], and YOLOX[45]. These models feature distinct network architectures, providing valuable insights into the performance of different object detection frameworks on denoised sonar images. Additionally, we employed the OPD[46] multi-frame image denoising algorithm to further denoise multiple denoised images, enhancing target visibility and reducing noise, thereby improving detection accuracy.
@@
Through comprehensive experiments, we conducted a comparative analysis of the differences in object detection results among the original dataset, the dataset with applied denoising algorithms, and the dataset with both denoising and multi-frame image denoising algorithms. This provides important reference for subsequent research in the field of underwater sonar image object detection.
We analyzed the effects of three advanced deep learning denoising algorithms on object detection in sonar images, offering valuable insights for future research in this domain.
We innovatively treated the denoised sonar images as multiple exposures of a single scene and input them into a multi-frame image denoising algorithm for integration. Experimental results validated the effectiveness of this approach.
II Related Work
II-A Object Detection
The methods for object detection in underwater sonar images can be categorized into traditional methods, CNN-based methods, and other machine learning methods. Traditional methods mainly rely on image processing and feature extraction techniques to accomplish object detection through manually designed rules and algorithms. Deep learning methods, particularly object detection techniques based on convolutional neural networks (CNN), have gradually demonstrated strong capabilities in sonar image processing. These methods learn features directly from data in an end-to-end manner, showing excellent adaptability and robustness. In addition to CNNs, traditional machine learning methods also play a role in object detection for underwater sonar images. These methods often combine handcrafted feature extraction with classifiers, offering flexibility but with performance limited by the quality of feature design.
Traditional Methods. Over the past decades, sonar image detection has predominantly relied on artificial feature extraction, mainly based on pixel values, features, and gray thresholds. Traditional underwater target detection methods typically depend on pixel value features, gray thresholds, or prior information regarding corresponding targets. For instance, [32] presents a hierarchical system that detects objects in underwater videos by transforming the input image into the spatial frequency domain and utilizing a frame-difference algorithm, demonstrating robust performance and accuracy across various environments. Another study [33] introduces a two-dimensional object segmentation method based on the OS-CFAR technique, designed to effectively extract targets from side-scan sonar data, showing improved resilience to noise and validating its effectiveness through experiments on real images. Additionally, a symbolic pattern analysis method [34] is proposed for robust feature extraction from sidescan sonar images, aimed at detecting underwater targets like mines, with real-time object recognition validated on actual naval operational data. Finally, another study [35] showcases preliminary results in change detection by surveying different seafloor environments, with the goal of developing an automated tool for detecting recently placed objects on the seabed.
CNN-Based Methods. In recent years, CNN-based methods have gained widespread adoption in underwater object detection. For instance, [24] employs CNN for object recognition in forward-looking sonar images, which significantly surpasses existing template matching methods, while requiring fewer learnable parameters to achieve high performance that generalizes well to unseen data. [15] introduces YOLOv3-DPFIN, an improved YOLOv3 [16] algorithm that employs a dual-path feature fusion network for real-time, noise-resistant underwater target detection. This method enhances feature extraction capabilities and multi-scale prediction accuracy, significantly boosting detection speed and precision. [38] proposes an improved YOLOv5-based target detection model for forward-looking sonar images, utilizing transfer learning and CoordConv to enhance feature extraction, resulting in improved detection accuracy and speed. Moreover, the detection network UFIDNet proposed in [18] integrates speckle noise reduction and scene prior knowledge to enhance detection performance, incorporating an auxiliary branch and feature selection strategy to optimize accuracy and efficiency. On the other hand, [14] presents EPL-UFLSID, a pseudo label-driven algorithm for underwater forward-looking sonar image object detection, which leverages high-quality pseudo labels to guide feature extraction for improved detection outcomes. Meanwhile, [19] introduces a semi-supervised framework based on a multibranch convolutional encoder-decoder network (MB-CEDN) for multitarget detection and segmentation in circular-scan synthetic aperture sonar (CSAS) images, significantly enhancing both detection and segmentation accuracy. Yang et al. [62] proposed a lightweight underwater object detection network, FLSD-Net, which enhances the detection accuracy and speed of small underwater objects by addressing overfitting and generalization issues, thereby providing greater robustness for practical applications.The following two studies explore the application of improved YOLOv7 models in underwater object detection. Qin et al. [63] introduced the YOLOv7C model, which incorporates an attention mechanism and the Multi-GnBlock module to enhance object detection performance in underwater sonar images with complex backgrounds. Additionally, pruning techniques were applied to improve real-time performance. Wen et al. [65], focusing on side-scan sonar images, integrated the Swin-Transformer and CBAM modules to strengthen the model’s attention to target regions. Furthermore, they innovatively introduced a feature scale factor, significantly improving detection accuracy.Both studies demonstrate the significant advantages of the improved YOLOv7 models in underwater object detection tasks.
Other Machine Learning Methods. In traditional methods for underwater object detection, [22] presents a cascade of boosted classifiers based on the Viola and Jones algorithm, emphasizing in-situ learning of target responses and local clutter. This approach enhances detection accuracy while minimizing false alarm rates in complex terrains, making it suitable for real-time applications. The method not only processes images rapidly but also adapts effectively to various complex environments. Additionally, reference [12] introduces an automatic detection method that combines clustering, segmentation, and feature extraction to achieve efficient and low-error detection without complex training processes.
II-B Denoising
Image denoising technology is crucial in the field of image processing, especially in noisy environments such as underwater sonar images. With the rapid development of deep learning, convolutional neural network-based methods have gradually become mainstream in the image denoising field. Deep learning methods offer significant advantages in denoising performance. For instance, RED-Net [48] proposed a deep fully convolutional network architecture with skip-layer connections to enhance image denoising performance. DnCNN [49] trains deep convolutional neural networks to effectively handle images with various noise intensities, making it particularly suitable for blind denoising tasks. FFDNet [50] introduces a noise level map, allowing the same network to handle images with varying noise levels, significantly improving inference speed. Additionally, CBDNet [51] adopts a more realistic noise model, combining synthetic and real noise images for training, effectively handling complex noise in real-world applications. In 2022, Chunwei Tian et al. [52] proposed MWDCNN, which combines convolutional neural networks with wavelet transform techniques and utilizes a multi-stage design to improve denoising performance. This method introduces dynamic convolution, allowing linear combinations between different convolution kernels to balance denoising performance and computational cost. In 2024, they proposed CTNet [53], which combines serial blocks, parallel blocks, and residual blocks, integrating the Transformer mechanism to further improve denoising performance. This model is more effective in handling complex noise patterns and is suitable for denoising tasks in large-scale datasets.
Unsupervised and self-supervised learning methods offer new ideas for image denoising, as they do not rely on clean images for training. For example, the Noise2Noise [54] method trains using pairs of noisy images, enabling effective recovery of clean images from noisy inputs. Noise2Self [55] and Noise2Void [56] utilize self-supervised learning, which does not require clean images or paired images. Instead, they perform denoising by exploiting the statistical properties of noise in the images. The Noisy-As-Clean [57] method treats noisy images as “clean” during training, effectively reducing the discrepancy between the noise and image priors, thereby improving denoising performance. Furthermore, Neighbor2Neighbor [2] and Blind2Unblind [5] are two recently proposed self-supervised denoising methods that can perform efficient denoising without the need for clean images.
For specific denoising issues in medical imaging and other domains, researchers have proposed plug-and-play denoising network modules. For example, LIDnet [58] integrates denoising tasks with lesion detection tasks and introduces an interest-region perception loss based on lesion detection, along with a collaborative training strategy, significantly improving denoising results and detection accuracy. Similarly, the DN-DETR [59] method incorporates a denoising training strategy to enhance the convergence speed and performance of the object detection method DETR.
In the field of underwater sonar image denoising, the UFIDNet [60] network proposed a method that combines speckle noise reduction auxiliary branches with feature selection strategies, significantly improving the accuracy and speed of underwater forward-looking sonar object detection. Meanwhile, Wang, X. et al. [61] proposed a denoising method based on the golden ratio for non-local spatial information and an adaptive cultural algorithm (NACA), showing better effectiveness and adaptability in both denoising and object detection tasks.
Overall, the denoising technology for underwater sonar images is still in the early stages of development. Future research needs to delve deeper into how to handle the various complex types of noise in underwater sonar images and better apply the mature denoising techniques developed for optical images to sonar images.
III Background
III-A Sonar Images
In seawater, due to changes in the medium, light waves and radio waves experience significant attenuation, resulting in very limited propagation distances. In contrast, sound waves propagate much better in water, covering a broader marine area. This makes sonar images more suitable for ocean exploration and various industries.
Sonar images can be classified into two types: side-scan sonar images and forward-looking sonar images. Side-scan sonar images are mainly used for exploring the seabed, generating images that present a lateral cross-section of the underwater terrain, typically used for target identification and underwater object exploration. On the other hand, forward-looking sonar images are used for scanning objects in front of the underwater vehicle, typically forming images in front of submarines or underwater robots, providing a more intuitive three-dimensional view to assist in identifying obstacles or targets ahead.

### v2

# Can Optical Denoising Clean Sonar Images? A Benchmark and Fusion Approach

@@ Abstract
Object detection in sonar images is crucial for underwater robotics applications including autonomous navigation and resource exploration. However, complex noise patterns inherent in sonar imagery, particularly speckle, reverberation, and non-Gaussian noise, significantly degrade detection accuracy. While denoising techniques have achieved remarkable success in optical imaging, their applicability to sonar data remains underexplored. This study presents the first systematic evaluation of nine state-of-the-art deep denoising models with distinct architectures, including Neighbor2Neighbor with varying noise parameters, Blind2Unblind with different noise configurations, and DSPNet, for sonar image preprocessing. We establish a rigorous benchmark using five publicly available sonar datasets and assess their impact on four representative detection algorithms: YOLOX, Faster R-CNN, SSD300, and SSDMobileNetV2. Our evaluation addresses three unresolved questions: first, how effectively optical denoising architectures transfer to sonar data; second, which model families perform best against sonar noise; and third, whether denoising truly improves detection accuracy in practical pipelines. Extensive experiments demonstrate that while denoising generally improves detection performance, effectiveness varies across methods due to their inherent biases toward specific noise types. To leverage complementary denoising effects, we propose a mutually-supervised multi-source denoising fusion framework where outputs from different denoisers mutually supervise each other at the pixel level, creating a synergistic framework that produces cleaner images. Experimental results show our proposed denoising approach boosts YOLOX’s detection performance by absolute margins of 22.8% (mAP@0.75) and 9.3% (mAP) when processing raw sonar images. The source code in this work is available at https://github.com/wzyii/M2SDF.
keywords:
[…]
1 Introduction
Underwater sonar imaging has become an indispensable tool for marine exploration and industrial applications, providing critical insights into subaquatic environments where optical systems face severe limitations [1, 2]. Unlike light waves that attenuate rapidly in water, acoustic signals propagate over significantly longer distances, enabling sonar systems to cover expansive underwater areas with practical resolution [3, 4, 5, 6]. This unique capability makes sonar imaging the preferred choice for seabed mapping, underwater infrastructure inspection, and autonomous underwater vehicle (AUV) navigation [7].
Despite these advantages, sonar image quality is frequently compromised by complex noise patterns that stem from three primary sources: 1) environmental interference (e.g., biological echoes and water column reflections), 2) sensor-related artifacts (e.g., transducer imperfections and electronic noise), and 3) inherent wave propagation characteristics (e.g., multipath reverberation and absorption loss) [8]. These manifest as distinct noise types including Gaussian noise from random sensor fluctuations, speckle noise caused by coherent wave interference, and Poisson noise due to signal quantum variability. Such degradations can reduce the signal-to-noise ratio (SNR) by over 15 dB in practical scenarios [9], severely impacting downstream tasks like mine detection or pipeline inspection where pixel-level accuracy is crucial.
Denoising presents a promising solution to enhance sonar image usability, but faces unique challenges compared to optical imaging: First, the physical mechanisms of acoustic wave scattering differ fundamentally from light scattering, requiring specialized treatment of multiplicative noise models. Second, the scarcity of annotated sonar datasets—typically orders of magnitude smaller than optical counterparts—constrains data-hungry deep learning approaches [10] [11] [12]. Third, the spatially varying nature of underwater noise distributions demands adaptive processing rather than uniform filtering [13].
Interestingly, modern deep learning-based denoisers developed for optical images demonstrate remarkable capability in handling similar challenges, particularly in preserving edges while suppressing noise [14]. The computational similarities between optical and sonar imaging (e.g., shared texture patterns and structural features) suggest potential for cross-domain adaptation. However, no systematic study has yet evaluated: 1) how effectively optical denoising architectures transfer to sonar data, 2) which model families perform best against sonar noise, or 3) whether denoising truly improves detection accuracy in practical pipelines.
To address these gaps, we conduct the first comprehensive evaluation of nine state-of-the-art denoising models across five public sonar datasets (FDD [15], UATD [16], SCTD [17], MarineDebris [18], and USD [19]). Our study reveals two key findings: 1) While denoising preprocessing improves detection mAP on average, performance gains vary substantially across methods; 2) No single denoiser achieves consistent superiority across all noise conditions, suggesting the need for adaptive solutions.
Building on these insights, we propose a mutually-supervised multi-source denoising fusion (M2SDF) framework that iteratively combines outputs from heterogeneous denoisers through pairwise pixel-level supervision. This learning mechanism identifies and preserves clean pixels while rectifying noise-contaminated regions, achieving sonar image enhancement.
@@ 1 Introduction
1.
We conduct the first large-scale evaluation of optical-based denoising methods for sonar imagery. Our benchmark validates the feasibility of transferring optical denoising techniques to sonar data and reveals fundamental limitations of single-model approaches, and provides practical guidelines for model selection based on noise characteristics.
2.
We present M2SDF, a novel multi-source denoising fusion framework that overcomes single-model limitations by combining pixel-wise mutual supervision with heterogeneous denoising strategies, improving sonar image denoising quality.
3.
The proposed M2SDF achieves significant performance gains in sonar image object detection using YOLOX, demonstrating absolute improvements of 22.8% in mAP@0.75 and 9.3% in mAP compared to noisy inputs.
2 Background
2.1 Sonar Images
In seawater, the propagation of light waves and radio waves is severely limited due to significant attenuation caused by medium variations. In contrast, sound waves exhibit superior propagation characteristics in aquatic environments, enabling broader coverage and making sonar imaging the preferred modality for ocean exploration and industrial applications. Sonar images can be categorized into two primary types: side-scan and forward-looking sonar images. Side-scan sonar images are primarily employed for seabed exploration, generating lateral cross-sectional representations of underwater terrain, which are instrumental in target identification and object detection. Forward-looking sonar images, on the other hand, are designed to scan the area ahead of underwater vehicles such as submarines or autonomous robots, providing an intuitive three-dimensional perspective to facilitate obstacle avoidance and target recognition. As shown in Fig.1, the top image depicts a side-scan sonar scene, and the bottom image shows a forward-looking sonar scene. In the side-scan image, a central black region indicates sonar shadows due to absent echoes beneath the detector, while high-intensity lateral regions reveal seabed topography, including geological structures like sand dunes.